Astrid1:

Calling upon for Math help: A cone has volume 36 pi. a. If the cones radius is 1, what is its height? b. If the cones radius is 2, what is its height? c. If the cones radius is 6, what is its height?

Shadow:

Let me just google the formula for the volume of a cone

Astrid1:

Okay!

Shadow:

\[V_{c} = \pi r \frac{ h }{ 3 }\]

Shadow:

It says the volume is 36 pi, which means we can say this: \[36 \pi = \pi(1)\frac{ h }{ 3 }\]

Shadow:

For part a

Shadow:

Since pi is on both sides, we can divide the entire equation by pi, turning it into \[36 = \frac{ h }{3 }\]

Shadow:

1 times h/3 = h/3, so I just simplified it there as well

Shadow:

Makes sense so far?

Astrid1:

Yeah

Shadow:

Do you know what you'd do next?

Astrid1:

Uhm, I am not sure.

Shadow:

In Algebra we want to isolate the variable, so we play with the numbers. However, in order to keep the equation true, whatever we do to one side, we do to the other. So what would get us h alone? Multiplying by 3 Since 3/3 = 1, and 1 times h is just h... \[3(36) = \frac{ h }{ 3 } \times 3 \rightarrow 3(36) = h \times 1 \rightarrow 3(36) = h\]

Shadow:

Whatever 3 times 36 is, is the height

Astrid1:

108?

Shadow:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow It says the volume is 36 pi, which means we can say this: \[36 \pi = \pi(1)\frac{ h }{ 3 }\] \(\color{#0cbb34}{\text{End of Quote}}\) One part I skipped over is that in the formula, it's actually r^2, so \[V_{cone} = \pi r^2 \frac{ h }{ 3 }\] But 1^2 is just 1 so I think my brain just kinda ignored it, but you'll need to respect the exponent for the other parts.

Shadow:

3 30s is 90, and 3 x 6 = 18, so yes

Astrid1:

Okay! Thanks.

Shadow:

Do you feel confident you can tackle the others?

Astrid1:

Yes, I do!

Shadow:

Cool

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow \[V_{c} = \pi r \frac{ h }{ 3 }\] \(\color{#0cbb34}{\text{End of Quote}}\) @Shadow excuse me but Volume formula of cone is area of base time height divide 3 so area of base will be pi*r^2 bc. is circle

jhonyy9:

@Astrid1 what is your question ?

Astrid1:

Nvm, Shadow helped me!

jhonyy9:

ok np yw